1. Is there a number that 1 compares to such that that number compares to itself plus 1? 1/x = x/x+1
Find me 2 solutions (non factorable)

Answer by stanbon(75887)

1/x = x/x+1
Cross-multiply and solve for "x":
x^2 = x+1
x^2-x-1 = 0
x = [1 +- sqrt(1-4*1*-1)]/2
---
x = [1 +- sqrt(5)]/2
--------------------------
x = (1+sqrt(5))/2 or x = (1-sqrt(5))/2
===========================================
Cheers,
Stan H.